A golf ball is driven horizontally from an elevated tee with a speed of 25 m/s. It strikes the fairway 2.5 sec later.

(a) How far has it fallen vertically?

(b) How far has it travelled horizontally?

(c) Find the horizontal and vertical components of its velocity just before it strikes the fairway. 


1
Expert's answer
2021-10-18T03:43:59-0400

"y = ut+ \frac12at²\\
y = 0(t) + \frac12(9.8)+(2.5)²\\
y = 0+ 4.9(6.25)\\
y = 30.625m"

(A)

Vertical distance will be : 30.625 m

(B)

Horizontal distance will be 62.5 m

(C) Components


Distance 


"d = \sqrt{x^2 + y^2 + z^2}"

Components of given forces 

"From \space \dfrac{F_x}{x} = \dfrac{F_y}{y} = \dfrac{F_z}{z} = \dfrac{F}{d}"

"F_x = \dfrac{x \, F_x}{d}"


Resultant

"R = \sqrt{{R_x}^2 + {R_y}^2 + {R_z}^2}"

Direction cosines of the resultant

"\cos \theta_x = \dfrac{R_x}{R} = -0.394"

"\cos \theta_y = \dfrac{R_y}{R} = 0.762"

"\cos \theta_z = \dfrac{R_z}{R} = -0.514"